three forces pull on object: f₁ = 50-N towards the +x; F₂ = 50-N toward the +y and F₃ = 70-N, 30° below -x. what force will balance the three? [Analythical method]
1
Expert's answer
2020-06-01T14:26:36-0400

Solution.

"F_1=50N;"

"F_2=50N;"

"F_3=70N;"

"\phi=30^o;"

"F-?;"


"\overrightarrow{F}=\overrightarrow{F_1}+\overrightarrow{F_2}+\overrightarrow{F_3};"

"\overrightarrow{F_{12}}=\overrightarrow{F_1}+\overrightarrow{F_2};"

Since the forces F1 and F2 are perpendicular, the equivalent of these forces is found by Pythagoras' theorem.

"F_{12}=\sqrt{F_1^2+F_2^2};"

"F_{12}=\sqrt{(50N)^2+(50N)^2}=50\sqrt{2}N;"

By the cosine theorem, we find the equivalent module:

"F^2=F_{12}^2+F_3^2-2F_{12}F_3cos\theta" ;

"F^2=(50\sqrt{2}N)^2+(70N)^2-2\sdot50\sqrt{2}N\sdot70N\sdot cos15^o="

"=339N^2;"

"F=18.4N;"

"\dfrac{F_3}{sin\alpha}=\dfrac{F}{sin\theta};\implies sin\alpha=\dfrac{F_3sin\theta}{F};"


"sin\alpha=\dfrac{70N\sdot 0.2588}{18.4N}=0.9845;"

"\alpha=80^o"

Equivalent force is directed at an angle of 55o above -x.

Answer: "F=18.4N" at an angle of 55o above -x.